SRPS DIESEL GROUP NEWS SECTION - UPDATED 02-05-2014
The metal frame work that was removed on the cab sides at number 2 end has now been installed with new metal, this has also allowed the cab sides themselves to begin, with B side cab side nearly completely welded with a new drop light drain channel being made and extensive repairs carried out around the quarter light. Refurbished exhausters have also been fitted back into the nose end. The body sides have seen progress where more poor metal has been removed with new sections being prepared for welding back into place. Shot blasting also continues on the body sides where required. Pictures of all the teams progress can be seen in the Smugmug gallery
The ETH cable fitting has now started at number 2 end, progressing through the loco towards number 1 end. New paths and routes are having to be found for the cable while ensuring the existing equipment is not influenced by the new cabling.
The loco is currently requiring an oil and filter change, it is currently out of traffic until this can be completed.
The shunter is now back in action after another repair to the water pump coupling.